An archer shoots an arrow at a 75.0 m distant target; the bull
Answer #1

a.)

I think that you can use the range equation to figure out the angle (?):

R=[(Vo)^2*2Sin?]/g

Where

R=73.0m
Vo=36m/s
g=9.8m/s^2
?=?

Solve for ?:

73.0m=[(36m/s)^2*2Sin?]/(9.8m/s^2)

73.0m*9.8m/s^2=[(36m/s)^2*2Sin?]

715.4=1296*2Sin?

(715.4/1296)*1/2=Sin?

(0.552)*1/2=Sin?

Sin^-1(0.276)=?

?=16.02